OpenStudy (anonymous):

WILL GIVE MEDAL TO BEST ANSWER w^12/w^18 =? 1. w^30 2. w^-30 3. w^6 4. w^-6

OpenStudy (anonymous):

@Kainui

OpenStudy (kainui):

Try thinking about a simple example like: x^3/x^2. If you write it out, you get (x*x*x)/(x*x) which simplifies to x. So you can see it's just 3-2=1 for the exponents.

OpenStudy (anonymous):

@Kainui So, it would be w^-6

OpenStudy (anonymous):

\[x^a * x^b = x^{a + b}\]\[\frac{ x^a }{ x^b } = x^{a-b}\]

OpenStudy (kainui):

Yeah you got it.

OpenStudy (anonymous):

Thanks!